OpenStudy (anonymous):

Someone please explain this to me. -A farmer has 400 feet of fencing. He wants to fence off a rectangular field that borders a straight river (no fence along the river). What are the dimensions that will give him the largest area?

OpenStudy (anonymous):

This is an optimization problem with a constraint. The farmer has 400 feet of fencing. That means that the total perimeter of the field is going to be 400 feet; or, letting l = length and w = width, 2L+2w=400. From that, we can solve for either L or w; in this case, we have w = 200 - L. What do we do with this? Well, we want to optimize area; area = Lw. Plugging in our substitution, A = L(200-L) = 200L-L^2. To optimize this with respect to L, we take the derivative with respect to L and set it equal to zero - 200 - 2L = 0 => L = 100. Thus a fence with a length of 100 feet will produce the optimal area. We can determine optimal width by plugging this into the perimeter equation, to get that the ideal dimensions are 100x100.

OpenStudy (anonymous):

but why does it turn into 200?

OpenStudy (anonymous):

Why does what turn into 200?

OpenStudy (anonymous):

oh just kidding i got it i got confused for a second thank you though!